# A statistically consistent measure of Semantic Variability using Language Models

### v2

@@ 5 Discussion
.. Many natural language processing tasks exhibit a fundamental invariance: sequences of distinct tokens can convey identical meanings. This paper introduces a theoretically grounded metric for quantifying semantic variation, referred to as semantic spectral clustering. This approach reframes the challenge of measuring semantic variation as a prompt-engineering problem, which can be applied to any large language model (LLM), as demonstrated through our simulation analysis. In addition, unsupervised uncertainty can offer a solution to the issue identified in prior research, where supervised uncertainty measures face challenges in handling distributional shifts.
While we define two texts as having equivalent meaning if and only if they mutually imply one another, alternative definitions may be appropriate for specific use cases. For example, legal documents could be clustered based on the adoption of similar legal strategies, with documents grouped together if they demonstrate comparable approaches. In such scenarios, the entropy of the legal documents could also be computed to quantify their informational diversity. We have demonstrated that, provided there exists a function $e$ capable of performing the evaluation with weak accuracy, this estimator remains consistent. Given the reasoning capabilities of large language models (LLMs), we foresee numerous possibilities for extending this method to a wide range of applications.
@@
6 Limitation
.. We acknowledge that, while this research offers a theoretically consistent measurement of variation, it does not account for situations where two pieces of text may partially agree. For instance, two texts may contain points of agreement as well as points of disagreement. This is particularly common when different authors cite the same sources but reach contradictory conclusions.
Appendix A Theoretical Result
A.1 Proof of proposition 2.1
@@ Appendix A Theoretical Result
Assumption A.1 implies that the spectral decomposition
@@
Theorem A.1.
(Collorary 2.2) Corollary 2.2. Suppose that Assumptions A.1, A.2, A.3, and A.4 hold and the $K$-means algorithm is applied to $\hat{\beta}_{in}=(n/K)^{1/2}\hat{u}_{1i}$ and $\beta_{g_{i}^{0}n}=(n/K)^{1/2}\hat{O}_{n}u_{1i}$ Then,
@@
Theorem A.4 (Theorem 3.2 with all conditions stated).
Assume that Assumptions A.1, A.2, A.3, and A.4 hold and the $K$-means algorithm is applied to $\hat{\beta}_{in}=(n/K)^{1/2}\hat{u}_{1i}$ and $\beta_{g_{i}^{0}n}=(n/K)^{1/2}\hat{O}_{n}u_{1i}$ Then
@@
Using Theorem A.1, we know that under Assumptions A.1, A.2, A.3, and A.4, we have that
@@
Using Lemma A.2, we know that
@@
A.3 Proof of Theorem 3.3
Theorem A.5 (Corollary 3.2 in Lei and Rinaldo (2015)).
@@
Under the model we have, we know that minimum eigenvalue of $B$ is $\lambda$. Use theorem A.5 to replace $h\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$ with $h\left(\frac{2K}{c_{2}}\right)c^{-1}(2+\varepsilon)\frac{Kn_{\max}}{n_{\min}^{2}\lambda^{2}\alpha_{n}}$ in lemma 3.1.
We now have to show the existence of $c$ in Theorem A.5.
@@
Appendix B Simulations
B.1 Hobby Examples
@@
Appendix C Prompt

### v3

@@ 5 Discussion
.. Many natural language processing tasks exhibit a fundamental invariance: sequences of distinct tokens can convey identical meanings. This paper introduces a theoretically grounded metric for quantifying semantic uncertainty, referred to as semantic spectral clustering. This approach reframes the challenge of measuring semantic uncertainty as a prompt-engineering problem, which can be applied to any large language model (LLM), as demonstrated through our simulation analysis. In addition, unsupervised uncertainty can offer a solution to the issue identified in prior research, where supervised uncertainty measures face challenges in handling distributional shifts. While we define two texts as having equivalent meaning if and only if they mutually imply one another, alternative definitions may be appropriate for specific use cases. For example, legal documents could be clustered based on the adoption of similar legal strategies, with documents grouped together if they demonstrate comparable approaches. In such scenarios, the entropy of the legal documents could also be computed to quantify their informational diversity. We have demonstrated that, provided there exists a function $e$ capable of performing the evaluation with weak accuracy, this estimator remains consistent. Given the reasoning capabilities of large language models (LLMs), we foresee numerous possibilities for extending this method to a wide range of applications.
@@
Appendix A Limitation
.. We acknowledge that, while this research offers a theoretically consistent measurement of variation, it does not account for situations where two pieces of text may partially agree. For example, two texts may contain points of agreement and points of disagreement. This is particularly common when different authors cite similar sources, but reach contradictory conclusions. Computing semantic similarity in this case is difficult.
Appendix B Theoretical Result
B.1 Proof of proposition 2.1
@@ Appendix B Theoretical Result
Assumption B.1 implies that the spectral decomposition
@@
Theorem B.1.
(Collorary 2.2) Corollary 2.2. Suppose that Assumptions B.1, B.2, B.3, and B.4 hold and the $K$-means algorithm is applied to $\hat{\beta}_{in}=(n/K)^{1/2}\hat{u}_{1i}$ and $\beta_{g_{i}^{0}n}=(n/K)^{1/2}\hat{O}_{n}u_{1i}$ Then,
@@
Theorem B.4 (Theorem 3.2 with all conditions stated).
Assume that Assumptions B.1, B.2, B.3, and B.4 hold and the $K$-means algorithm is applied to $\hat{\beta}_{in}=(n/K)^{1/2}\hat{u}_{1i}$ and $\beta_{g_{i}^{0}n}=(n/K)^{1/2}\hat{O}_{n}u_{1i}$ Then
@@
Using Theorem B.1, we know that under Assumptions B.1, B.2, B.3, and B.4, we have that
@@
Using Lemma B.2, we know that
@@
B.3 Proof of Theorem 3.3
Theorem B.5 (Corollary 3.2 in Lei and Rinaldo (2015)).
@@
Under the model we have, we know that minimum eigenvalue of $B$ is $\lambda$. Use theorem B.5 to replace $h\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$ with $h\left(\frac{2K}{c_{2}}\right)c^{-1}(2+\varepsilon)\frac{Kn_{\max}}{n_{\min}^{2}\lambda^{2}\alpha_{n}}$ in lemma 3.1.
We now have to show the existence of $c$ in Theorem B.5.
@@
Appendix C Simulations
C.1 Hobby Examples
@@
C.2 Historical Examples
On the day December 3,
  • "Pope John X crowns Berengar I of Italy as Holy Roman Emperor"/ "Berengar I of Italy was crowned Emperor by Pope John X."
  • "USS Alfred becomes the first vessel to fly the Grand Union Flag; the flag is hoisted by John Paul Jones."/ "John Paul Jones hoisted the the Grand Union Flag on USS Alfred, the first vessel to fly Grand Union Flag."
  • "French General Jean Victor Marie Moreau decisively defeats the Archduke John of Austria near Munich in the Battle of Hohenlinden. "/"Archduke John of Austria was defeated by French General Jean Victor Marie Moreau near Munich in the Battle of Hohenlinden."
  • "Illinois becomes the 21st U.S. state."/ "Illinois joined U.S. as its 21st state."
  • "The Zollverein (German Customs Union) begins the first regular census in Germany"/"The first regular census was conducted by The Zollverein (German Customs Union) in Germany."
  • "The Duquesne Country and Athletic Club defeats an all-star collection of early football players 16–0."
  • "Following more than a month of Turkish–Armenian War, the Turkish-dictated Treaty of Alexandropol is concluded."/ "The Turkish-dictated Treaty of Alexandropol concluded after a month of Turkish–Armenian War."
  • "President Herbert Hoover delivers his first State of the Union message to Congress. It is presented in the form of a written message rather than a speech."/"President Herbert Hoover presents a written message as his first State of the Union message to Congress rather than a speech."
  • "The current flag of Singapore is adopted, six months after Singapore became self-governing within the British Empire."/"Singapore adopts it current flag six months after it self-govern within the British Empire."
  • "Ayatollah Ruhollah Khomeini becomes the first Supreme Leader of Iran.", "Iran has its first Supreme Leader: Ayatollah Ruhollah Khomeini."
Appendix D Prompt
@@
Appendix E Notes on Implementation in python
The default spectral clustering algorithm in python already implemented in k++ algorithm Choo et al. (2020).